# for the Hadron Spectrum Collaboration

### v1

@@ I Introduction
The wealth of exotic hadron states discovered recently in experimentation has driven theoretical efforts to understand their nature. In particular, the recent observation of the doubly-charmed, manifestly exotic $T_{cc}^{+}(3875)$ at the LHCb experiment [1, 2] close in energy to ${DD^{\ast}}$ threshold motivates further study. Having two-charm quarks and being charged, this state goes beyond expectations from conventional quark model $q\bar{q}$ spectra, with a minimal $cc\bar{u}\bar{d}$ quark content content [1]. LHCb investigates several charged and neutral ${DD^{\ast}}$ combinations, finding only one enhancement and implying a preference for isospin $I=0$.
@@
Quantum Chromodynamics (QCD) is the theory of the strong interactions and it is thus desirable to know what QCD predicts for the $T_{cc}$. A nonperturbative approach to QCD making only controlled and systematically improvable approximations is lattice QCD, where the fundamental equations of interacting quarks and gluons are discretized on a finite spacetime lattice. The QCD path integral is sampled on the lattice using Monte Carlo techniques, which allows for the extraction of correlation functions and the calculation of the spectrum of the eigenstates of QCD on the lattice. Due to the finite spatial volume of the lattice, momentum is quantized and only a discrete spectrum is obtained.
The discrete spectra can then be used to determine infinite volume scattering amplitudes that are continuous functions of energy using the Lüscher finite volume formalism [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
The scattering amplitudes contain singularities, which are a model independent way of understanding the spectroscopic content in terms of bound states, virtual bound states, and resonances. The approach is reviewed in Ref. [28].
@@
Investigations of ${DD^{\ast}}$ on the lattice have been made over a range of light quark masses. Ref. [39] found that the ground state energy obtained utilizing a basis of meson-meson and tetraquark operators in $J^{P}=1^{+}$ lies below the ${DD^{\ast}}$ threshold using three lattice spacings and $m_{\pi}$ ranging from 153 to 689 MeV. This is suggestive of the presence of attractive interactions, however the existence of a pole was not verified. In Ref. [40], the HAL QCD method was used to extract $S$-wave phase shifts with pions at nearly physical mass, locating a virtual bound state 45 keV below ${DD^{\ast}}$ threshold. Ref. [41] examined elastic ${DD^{\ast}}$ scattering in $S$-wave for $I=0,1$, but did not verify the presence of a pole. Elastic ${DD^{\ast}}$ $S$-wave scattering was investigated in [42] for $m_{\pi}\approx 280$ MeV, and a pole in the scattering amplitudes corresponding to a virtual bound state was found $9_{-7}^{+4}$ MeV below the ${DD^{\ast}}$ threshold. Its quark mass dependence was examined in [43].
@@ II Partial waves
We first consider the possible partial waves and their relevant symmetries for ${DD^{\ast}}$ and ${D^{\ast}{D}^{\ast}}$ with exact isospin symmetry. These are classified by their total spin parity $J^{P}$, but for any given $J^{P}$ different orbital angular momenta may contribute due to the presence of the vector $D^{\ast}$. It is convenient to use a spectroscopic notation ${\>\!}^{{2S+1}\!}{\ell}_{J}$ with total spin $S$ and orbital angular momentum $\ell$. Partial waves grow at threshold like $k_{cm}^{2\ell}$ for cm momentum $k_{cm}$ in the absence of any enhancement.
@@ III.1 Lattice Setup
Lattice calculations were performed on anisotropic lattices with three volumes having the same quark mass and lattice spacing. The volumes and the number of configurations per volume are given in Table 3. In order to improve energy resolution, the temporal lattice spacing $a_{t}$ is chosen to be finer than the spatial one $a_{s}$ by a factor $\xi=a_{s}/a_{t}\approx 3.5$. Gauge configurations were generated using a tree level improved Symanzik action in the gauge sector and a Wilson-Clover action with $2+1$ flavors of dynamical quarks in the fermion sector, where the strange quark has been tuned to approximately its physical value [44, 45]. The mass of the $\Omega$ baryon is used to set the scale when calculating quantities in physical units, which is $a_{t}^{-1}=m^{\Omega}_{phys}/a_{t}m_{\Omega}=5667$ MeV [46], giving a pion mass of $m_{\pi}\approx 391$ MeV. This corresponds to a spatial lattice spacing $a_{s}\approx 0.12$ fm. The quenched charm quark also uses the Wilson-Clover action and its mass is tuned to approximately reproduce the physical mass of the $\eta_{c}$ meson [47]. This results in stable $D$ and $D^{\ast}$ mesons with masses approximately $1886$ MeV and $2010$ MeV respectively. In the scattering analyses, we use the anisotropy determined from the dispersion relation of the pion as the central value, which was found to be $\xi=\xi_{\pi}=3.444(6)$ [48]. However, to account for the systematic uncertainty from obtaining different anisotropies from the relativistic dispersion relations for the $\pi$, $D$ and $D^{\ast}$ mesons across the three lattice volumes, we utilize a value of $\xi=3.444(50)$ in this analysis. This uncertainty spans the central values of the anisotropies determined from the $\pi$, $D$ and $D^{\ast}$ mesons across the three lattice volumes and the uncertainity is propagated through to the cm-frame energies obtained from the lattice energies. Consequently, this is also propagated through into the scattering amplitudes. See [33, 38] for further details of the dispersion relations.
@@
Table 3: A summary of the details for each ensemble used in this work. $(L/a_{s})^{3}\times(T/a_{t})$ denotes the volume of the lattice with $L$ the spatial extent and $T$ the temporal extent of the lattice. $N_{\mathrm{cfgs}}$ denotes the number of gauge-field configurations used. $N_{\mathrm{tsrcs}}$ denotes the number of time sources used per gauge-field configuration. $N_{\mathrm{vecs}}$ denotes the number of vectors used in the distillation basis.
@@
Table 4: (Left) The relevant stable masses for the $\pi$ [48], $D$, $D^{\ast}$ [38] and $D_{0}^{\ast}$ [31] and (right) kinematic thresholds for this study.
@@ III.2 Details of the Calculation
Because the finite spatial volume and discretization of the lattice break the rotational symmetry of the the infinite volume continuum, angular momentum $J$ is no longer a good quantum number, and mesons at rest with definite continuum $J$ are labelled by the finite cubic group irreducible representation (irrep) $\Lambda$, which we refer to as subduction [53, 52]. Additionally, the rotational symmetry is further broken for mesons in flight, and so such states are labelled by an irrep of the little group LG$(\vec{P})$ [54, 55]. The subductions for pseduoscalar-vector and vector-vector scattering relevant to $D^{(\ast)}D^{\ast}$ can be found in Appendix A.
The basis of interpolating operators $\{\mathcal{O}_{i}\}$ must be chosen to be sufficiently large and with suitable structures such that all levels in the energy region of interest can be extracted reliably.
We employ meson-meson-like operators [48] of the form
@@
At the light-quark masses chosen, the first inelastic three-hadron channel is $DD\pi$, given in Table 4. This leads to a large window in energy where only ${DD^{\ast}}$ and ${D^{\ast}{D}^{\ast}}$ are kinematically open. Included in the basis are all two hadron operators which in the absence of interactions correspond to a two hadron cm-frame energy, i.e. non-interacting energy, below the $DD\pi$ threshold. Also included are a limited amount of additional operators whose non interacting energies lie above the $DD\pi$ threshold to ensure a robust basis. In [57], it was shown that the inclusion of local compact $qq\bar{c}\bar{c}$ operators had no significant effect on the extracted finite volume spectra of ${DD^{\ast}}$ and ${D^{\ast}{D}^{\ast}}$-like states in $I=0$, $J^{P}=1^{+}$ on the $16^{3}$ lattice used here, so they are not included in the basis. The complete bases of operators used to interpolate the $D^{(\ast)}D^{\ast}$ states, both for zero and nonzero total momentum, can be found in Appendix B.
Distillation vectors [58] are utilized in order to obtain the necessary two-point functions to form the correlation matrix of Eq. (1). The distillation basis is formed using the lowest eigenmodes of the 3D gauge covariant Laplacian restricted to a time-slice $t$ of the lattice, which projects out the high frequency modes of the gauge fields. This projection creates smeared quark fields that are used in the interpolating operators, allowing reliable extraction of many states in the finite volume spectrum, such as those resembling mesons with various lattice-permitted momenta. Additionally, the use of distillation vectors allows for the required Wick contractions to be efficiently computed. The number of distillation vectors and the number of time sources are given in Table 3.
@@ III.3 Finite Volume Spectra
Here the finite volume spectra in the $I=0$ doubly-charmed sector are presented. The irreps are labelled $[n_{x}n_{y}n_{z}]\Lambda^{(P)}$, where $[n_{x}n_{y}n_{z}]=\frac{2\pi}{L}(n_{x},n_{y},n_{z})=\vec{P}$ is the quantized momentum, while $\Lambda$ labels the irrep with parity $P$. For $\vec{P}=0$, we calculate the spectrum for the $T_{1}^{+},A_{1}^{-},E^{-},E^{+}$ and $A_{2}^{+}$ irreps, using the basis of operators given in Table 11 of Appendix B. Also calculated is the spectra for some irreps with $\vec{P}\neq 0$, namely $[001],[002],[011]$ and $[111]A_{2}$, using the basis of operators given in Table 12 of Appendix B. The use of the $A_{2}$ irreps allows us to constrain the $J^{P}=1^{+}$ scattering amplitudes in the full coupled-channel region, where the $[000]T^{+}_{1}$ irrep would offer only a small amount of constraint. The additional rest frame irreps $A_{1}^{-},E^{-},E^{+}$ and $A_{2}^{+}$ allow us to constrain other partial waves that subduce into $[000]T_{1}^{+}$ and the $A_{2}$ irreps. In Tables 7, 8, 9 and 10 in Appendix A we display the relevant partial wave subductions into these irreps for pseudoscalar-vector [56] and identical vector-vector meson scattering.
Figures 1 and 2 show the extracted finite volume spectra for the irreps with $\vec{P}=0$ and $\vec{P}\neq 0$, respectively. The relevant kinematic thresholds and non interacting energies, the expected cm-frame energies in the absence of any interaction, are labeled with their degeneracy. Points in black are those used in the subsequent scattering analysis, while those in grey are not used. In the $[000]T_{1}^{+}$ and $[002]A_{2}$ irreps, the ground state energies display significant shifts down from the ${DD^{\ast}}$ threshold, which is similar to previous lattice studies in $I=0$. In addition, the energy levels near the ${D^{\ast}{D}^{\ast}}$ threshold have significant shifts downwards, consistent with the spectrum obtained in [57] on a single lattice volume. A similar picture is present in the $[n_{x}n_{y}n_{z}]A_{2}$ irreps. While the degeneracies of the non interacting levels are large in these irreps, we observe that some energy levels near ${D^{\ast}{D}^{\ast}}$ threshold have significant shifts down away from the non interacting levels.
@@
As discussed in Ref. [38], there is a tension between the $D$ and $D^{\ast}$ meson masses on the three volumes due to some unaccounted for uncertainty. We allow for this by adding an additional systematic uncertainty to the finite-volume energy levels using
@@
where $a_{t}\delta E_{\mathrm{latt}}$ is the uncertainty of the energy level and we set $a_{t}\delta E_{\mathrm{syst}}=5\times 10^{-4}$. The addition of this systematic error relieves the tension in the observed masses obtained from the relativistic dispersion fits of the $D$ and $D^{\ast}$ which may have been attributed to discretization effects or other additional sources of unknown systematic error. This added systematic error is propagated through to the uncertainty in the final scattering amplitudes. Further details can be found in [38] in the context of charmonium scattering and [33] in the context of $DK$ scattering.

### v2

@@ I Introduction
The wealth of exotic hadron states discovered recently in experimentation has driven theoretical efforts to understand their nature. In particular, the recent observation of the doubly-charmed, manifestly exotic $T_{cc}^{+}(3875)$ at the LHCb experiment [1, 2] close in energy to ${DD^{\ast}}$ threshold motivates further study. Having two charm quarks and being charged, this state goes beyond expectations from conventional quark model $q\bar{q}$ spectra, with a minimal $cc\bar{u}\bar{d}$ quark content content [1]. LHCb investigates several charged and neutral ${DD^{\ast}}$ combinations, finding only one enhancement and implying a preference for isospin $I=0$.
@@
Quantum Chromodynamics (QCD) is the theory of the strong interactions and it is thus desirable to know what QCD predicts for the $T_{cc}$. A nonperturbative approach to QCD making only controlled and systematically improvable approximations is lattice QCD, where the fundamental equations of interacting quarks and gluons are discretized on a finite spacetime lattice. The QCD path integral is sampled on the lattice using Monte Carlo techniques, which allows for the extraction of correlation functions and the calculation of the spectrum of the eigenstates of QCD on the lattice. Lattice calculations are performed in a finite spatial volume and we use periodic spatial boundary conditions. As such, the momentum is quantized and a discrete spectrum is obtained.
The discrete spectra can then be used to determine infinite volume scattering amplitudes that are continuous functions of energy using the Lüscher finite volume formalism [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. The scattering amplitudes contain singularities, which are a model independent way of understanding the spectroscopic content in terms of bound states, virtual bound states, and resonances. The approach is reviewed in Ref. [28].
@@
Investigations of ${DD^{\ast}}$ on the lattice have been made over a range of light quark masses. Ref. [39] found that the ground state energy obtained utilizing a basis of meson-meson and tetraquark operators in $J^{P}=1^{+}$ lies below the ${DD^{\ast}}$ threshold using three lattice spacings and $m_{\pi}$ ranging from 153 to 689 MeV. This is suggestive of the presence of attractive interactions, however the existence of a pole was not verified. In Ref. [40], the HAL QCD method was used to extract $S$-wave phase shifts with pions at nearly physical mass, locating a virtual bound state 45 keV below ${DD^{\ast}}$ threshold. Ref. [41] examined elastic ${DD^{\ast}}$ scattering in $S$-wave for $I=0,1$, but did not verify the presence of a pole. Elastic ${DD^{\ast}}$ $S$-wave scattering was investigated in [42] for $m_{\pi}\approx 280$ MeV, and a pole in the scattering amplitudes corresponding to a virtual bound state was found $9_{-7}^{+4}$ MeV below the ${DD^{\ast}}$ threshold. Its quark mass dependence and the role of the left hand cut due to one pion exchange between $D$ and $D^{\ast}$ were examined in [43]. A left hand cut due to one pion exchange between $D$ and $D^{\ast}$ is also present in this calculation, as well as a second left hand cut due to one pion exchange between $D^{\ast}$ and $D^{\ast}$ in the coupled-channel case.
@@ II Partial waves
We first consider the possible partial waves and their relevant symmetries for ${DD^{\ast}}$ and ${D^{\ast}{D}^{\ast}}$ with exact isospin symmetry. These are classified by their total spin parity $J^{P}$, but for any given $J^{P}$ different orbital angular momenta may contribute due to the presence of the vector $D^{\ast}$. It is convenient to use a spectroscopic notation ${\>\!}^{{2S+1}\!}{\ell}_{J}$ with total spin $S$ and orbital angular momentum $\ell$. Partial waves grow at threshold like $k_{cm}^{2\ell}$ in the absence of any enhancement, where $k_{cm}=\sqrt{(s-(m_{D^{(\ast)}}+m_{D^{\ast}})^{2})(s-(m_{D^{(\ast)}}-m_{D^{\ast}})^{2})/4s}$ is the scattering momentum in the center-of-mass (cm) frame and $s=E_{cm}^{2}$ with $E_{cm}$ the total energy in the cm-frame.
@@ III.1 Lattice Setup
Lattice calculations were performed on anisotropic lattices with three volumes having the same quark mass and lattice spacing. The volumes and the number of configurations per volume are given in Table 3. In order to improve energy resolution, the temporal lattice spacing $a_{t}$ is chosen to be finer than the spatial one $a_{s}$ by a factor $\xi=a_{s}/a_{t}\approx 3.5$. Gauge configurations were generated using a tree level improved Symanzik action in the gauge sector and a Wilson-Clover action with $2+1$ flavors of dynamical quarks in the fermion sector, where the strange quark has been tuned to approximately its physical value [44, 45]. The mass of the $\Omega$ baryon is used to set the scale when calculating quantities in physical units, which is $a_{t}^{-1}=m^{\Omega}_{phys}/a_{t}m_{\Omega}=5667$ MeV [46], giving a pion mass of $m_{\pi}\approx 391$ MeV. This corresponds to a spatial lattice spacing $a_{s}\approx 0.12$ fm. The quenched charm quark also uses the Wilson-Clover action and its mass is tuned to approximately reproduce the physical mass of the $\eta_{c}$ meson as discussed in [47], which results in a bare mass of $a_{t}m_{c}=0.092$. This results in stable $D$ and $D^{\ast}$ mesons with masses approximately $1886$ MeV and $2010$ MeV respectively. See Table 4 for the masses of the relevant stable mesons in this study. In the scattering analyses, we use the anisotropy determined from the dispersion relation of the pion as the central value, which was found to be $\xi=\xi_{\pi}=3.444(6)$ [48]. However, to account for the systematic uncertainty from obtaining different anisotropies from the relativistic dispersion relations for the $\pi$, $D$ and $D^{\ast}$ mesons across the three lattice volumes, we utilize a value of $\xi=3.444(50)$ in this analysis. This uncertainty spans the central values of the anisotropies determined from the $\pi$, $D$ and $D^{\ast}$ mesons across the three lattice volumes and the uncertainty is propagated through to the cm-frame energies obtained from the lattice energies. The lattice energies are calculated in the lattice frame where the $D^{(\ast)}D^{\ast}$ system can have overall nonzero momentum and are translated to the cm-frame where it has overall zero momentum. Consequently, this is also propagated through into the scattering amplitudes. See [33, 38] for further details of the dispersion relations.
@@
Table 3: A summary of the details for each ensemble used in this work. $(L/a_{s})^{3}\times(T/a_{t})$ denotes the volume of the lattice with $L$ the spatial extent and $T$ the temporal extent of the lattice. $N_{\mathrm{cfgs}}$ denotes the number of gauge-field configurations used. $N_{\mathrm{tsrcs}}$ denotes the number of time sources used per gauge-field configuration. On the $(L/a_{s})^{3}=20^{3},24^{3}$ volumes, four time sources were used to calculate the spectrum for overall zero momentum and three for overall nonzero momentum.
$N_{\mathrm{vecs}}$ denotes the number of vectors used in the distillation basis.
@@
Table 4: (Left) The relevant masses for the stable $\pi$ [48], $D$, $D^{\ast}$ [38] and $D_{0}^{\ast}$ [31] and (right) kinematic thresholds for this study.
@@ III.2 Details of the Calculation
Because the finite spatial volume and discretization of the lattice break the rotational symmetry of the infinite volume continuum, angular momentum $J$ is no longer a good quantum number, and mesons at rest with definite continuum $J$ are labelled by the finite cubic group irreducible representation (irrep) $\Lambda$, which we refer to as subduction [53, 52]. Additionally, the rotational symmetry is further broken for mesons in flight, and so such states are labelled by an irrep of the little group LG$(\vec{P})$ [54, 55]. The subductions for pseduoscalar-vector and vector-vector scattering relevant to $D^{(\ast)}D^{\ast}$ can be found in Appendix A.
The basis of interpolating operators $\{\mathcal{O}_{i}\}$ must be chosen to be sufficiently large and with suitable structures such that all levels in the energy region of interest can be extracted reliably. We employ meson-meson-like operators [48] of the form
@@
At the light-quark masses chosen, the first three-hadron channel is $DD\pi$, given in Table 4. This leads to a large window in energy where only ${DD^{\ast}}$ and ${D^{\ast}{D}^{\ast}}$ are kinematically open. Included in the basis are all two hadron operators which in the absence of interactions correspond to a two hadron cm-frame energy, i.e. non-interacting energy, below the $DD\pi$ threshold. Also included are a limited amount of additional operators whose non interacting energies lie above the $DD\pi$ threshold to ensure a robust basis. In [57], it was shown that the energy levels obtained using meson-meson operators were consistent within statistical uncertainty to the energy levels obtained using both meson-meson and local compact $cc\bar{q}\bar{q}$ operators in the doubly charmed sector with $I=0,J^{P}=1^{+}$ (see Figures 6 and 8 within Ref. [57] ). [^1] Ref. [57] did not relate the finite-volume energy levels to scattering amplitudes because only one relatively small volume was used and not enough energy levels were extracted to reliably determine the scattering amplitudes. As such, we do not include local compact $cc\bar{q}\bar{q}$ operators in this study. The complete bases of operators used to interpolate the $D^{(\ast)}D^{\ast}$ states, both for zero and nonzero total momentum, can be found in Appendix B.
Distillation vectors [58] are utilized in order to obtain the necessary two-point functions to form the correlation matrix of Eq. (1). The distillation basis is formed using the lowest eigenmodes of the 3D gauge covariant Laplacian restricted to a time-slice $t$ of the lattice, which projects out the high frequency modes of the gauge fields. This projection creates smeared quark fields that are used in the interpolating operators, allowing reliable extraction of many states in the finite volume spectrum, such as those resembling mesons with various lattice-permitted momenta. Additionally, the use of distillation vectors allows for the required Wick contractions to be efficiently computed. The number of distillation vectors and the number of time sources are given in Table 3. The four irreps with overall nonzero momentum used three time sources. In order to reduce correlations, each non-zero momentum irrep used a different combination of three time sources from the four available.
@@ III.3 Finite Volume Spectra
Here the finite volume spectra in the $I=0$ doubly-charmed sector are presented. The irreps are labelled $[n_{x}n_{y}n_{z}]\Lambda^{(P)}$, where $[n_{x}n_{y}n_{z}]=\frac{2\pi}{L}(n_{x},n_{y},n_{z})=\vec{P}$ is the quantized momentum, while $\Lambda$ labels the irrep with parity $P$. For $\vec{P}=0$, we calculate the spectrum for the $T_{1}^{+},A_{1}^{-},E^{-},E^{+}$ and $A_{2}^{+}$ irreps, using the basis of operators given in Table 11 of Appendix B. Also calculated are the spectra for some irreps with $\vec{P}\neq 0$, namely $[001],[002],[011]$ and $[111]A_{2}$, using the basis of operators given in Table 12 of Appendix B. The use of the $A_{2}$ irreps allows us to constrain the $J^{P}=1^{+}$ scattering amplitudes in the full coupled-channel region, where the $[000]T^{+}_{1}$ irrep would offer only a small amount of constraint. The additional rest frame irreps $A_{1}^{-},E^{-},E^{+}$ and $A_{2}^{+}$ allow us to constrain other partial waves that subduce into $[000]T_{1}^{+}$ and the $A_{2}$ irreps. In Tables 7, 8, 9 and 10 in Appendix A we display the relevant partial wave subductions into these irreps for pseudoscalar-vector [56] and identical vector-vector meson scattering.
Figures 1 and 2 show the extracted finite volume spectra for the irreps with $\vec{P}=0$ and $\vec{P}\neq 0$, respectively. The relevant kinematic thresholds and non interacting energies, the expected cm-frame energies in the absence of any interaction, are labeled with their degeneracy. Points in black are those used in the subsequent scattering analysis, while those in grey are not used. In the $[000]T_{1}^{+}$ and $[002]A_{2}$ irreps, the ground state energies display significant shifts down from the ${DD^{\ast}}$ threshold, which is similar to previous lattice studies in $I=0$. In addition, the energy levels near the ${D^{\ast}{D}^{\ast}}$ threshold have significant shifts downwards. The spectrum obtained on the $(L/a_{s})^{3}=16^{3}$ lattice volume for $[000]T_{1}^{+}$ is consistent with the spectrum obtained in [57] both with and without local compact $cc\bar{q}\bar{q}$ operators. A similar picture is present in the $[n_{x}n_{y}n_{z}]A_{2}$ irreps. While the degeneracies of the non interacting levels are large in these irreps, we observe that some energy levels near ${D^{\ast}{D}^{\ast}}$ threshold have significant shifts down away from the non interacting levels. We remark that all energy levels found in this calculation lie above the opening of the left hand cut due to one pion exchange between the $D$ and $D^{\ast}$, which begins at $a_{t}E_{cm}=0.68432$. Another left hand cut due to one pion exchange between $D^{\ast}$ and $D^{\ast}$ begins at $a_{t}E_{cm}=0.70591$. We will return to the discussion of the two left hand cuts in Section VI.1.
@@
As discussed in Ref. [38], there is a tension between the $D$ and $D^{\ast}$ meson masses on the three volumes due to some unaccounted for uncertainty (see Figure 25 of [38]). We allow for this by adding an additional systematic uncertainty to the finite-volume energy levels using
@@
where $a_{t}\delta E_{\mathrm{latt}}$ is the uncertainty of the energy level and we set $a_{t}\delta E_{\mathrm{syst}}=5\times 10^{-4}$ as used in [38]. The addition of this systematic error relieves the tension in the observed masses obtained from the relativistic dispersion fits of the $D$ and $D^{\ast}$ which may have been attributed to discretization effects or other additional sources of unknown systematic error. This added systematic error is propagated through to the uncertainty in the final scattering amplitudes. Further details can be found in Appendix A of [38] in the context of charmonium scattering and [33] in the context of $DK$ scattering. This tension has been observed in other lattice calculations using charm quarks [36, 59, 60] and various methods to relieve this tension can be employed. We view our approach as a relatively conservative method.